4x+13=-9-3(x+9);x=-7
Answer by BrittanyM(80)

This problem is asking for us to prove that each side of the equation is equal when x is equal to negetive seven. So we plug negetive seven into the equation in place of x.
4(-7)+13=-9-3[(-7)+9]
-28+13=-9-3(2)
-15=-9-6
-15=-15